1. Consider a generator that rotates its 200 turn, 0.13 m diameter coil at 3550 rpm in a 0.65 T field.

a. Calculate the peak voltage of the generator.

2. A precision laboratory resistor is made of a coil of wire. The coil is 1.5 cm in diameter, 3.75 cm long, and has 500 turns.

a.

What is its self-inductance in mH?

b. What average emf in V is induced if the 12.0 A current through it is turned on in 5.00 ms (one-fourth of a cycle for 50 Hz AC)?  
Answer #1

1)Peak voltage=BAN\omega

B=magnetic field=.65T

A=area of cross section of coil=77 =3.14X.065^{2}=.01327m2(r=radius=.13/2=.065m

N=number of turns=200

\omega=angular velocity=2\pif

f=frequency of rotation=3550rpm=3550/60=59.17Hz

\omega=2X3.14X59.17=371.59rad/s

Peak voltage=BAN\omega=.65X.01327X200X371.59=641.03volt

2

a)Self inductance L=HON^{2}A/l

HO=4\piX10:

N=number of turns=500

A=area of cross section=\pir^{2}=3.14X.0075^{2}=1.77X10^{-4}m2 (r=1.5/2=.75cm)

l=length=3.75cm=.0375m

L=HON^{2}A/l=4\piX10:x500^{2}X1.77X10^{-4}/.0375=1.48mH

b)induced emf e=Ldi/dt

di=change in current=12-0=12A

dt=time=5ms=.005

L=1.48mH=.00148H

e=Ldi/dt=.00148x12/.005=3.552volt